Exercise 13A-3 Basic Present Value Concepts [LO13-7]

In three years, when he is discharged from the Air Force, Steve wants to buy an $13,000 power boat.

Click here to view Exhibit 13B-1 and Exhibit 13B-2, to determine the appropriate discount factor(s) using tables.

Required:

What lump-sum amount must Steve invest now to have the $13,000 at the end of three years if he can invest money at: (Round your final answer to the nearest whole dollar amount.)
Answer #1

Use this below table to calculate Present Value of desired amount: -

3 А 1 Future Value 2 Time period 3 Discounting Rate =10/100 4 Discounting Factor =1/(1+B3) B2 5 Present Value =B1*B4 Assumed

End results would depend upon your discounting rate or rate which is provided: -

А 1 Future Value 13000 2 Time period 3 Discounting Rate 0.11 Assumed 4 Discounting Factor 0.751314801 5 Present Value 9767.09